Use the equations below to answer the following questions


P=60-1/3Qd


2p=40+2/5Qs


Calculate the equilibrium price and quantity


If the is set at p=40,state whether it is a price ceiling or price floor and calculate the size of imbalance


Calculate the price elasticity of demand and supply

1
Expert's answer
2022-02-18T15:51:47-0500

Equilibrium Price and Quantity.

"P=60-\frac{1}{3}Q_d"

"Q_d=180-3P"

"2P=40+\frac{2}{5}Q_s"

"Q_s=5P-100"

at equilibrium, "Q_d=Q_s"

"\therefore 180-3P=5P-100"

"280=8P"

"\bold{P=35}"

Replacing the value of P in either the "Q_d" or "Q_S" we get the value of quantity as;

"180-3(35)=75"

"\bold{Q=75}"


When the price is set at "P=40" , it is to be considered as a price ceiling. Since the price ceiling is above the equilibrium price, then there is no imbalances arising in the market.


Price Elasticity(Ped)

"Ped=\frac{\delta Q/Q}{\delta P/P}"

at "P=40," "Q=180-3(40)=60"

"\delta Q=60-75=-15"

"\delta P= 40-35=5"

"Ped=\frac{-15/60}{5/40}=-2"

"\bold{Ped=-2}"